OpenStudy (anonymous):

Evaluate f-2g for f = 4 and g = 32.

OpenStudy (anonymous):

(4)-2(32)=4-64=-60

Nnesha (nnesha):

just plug g and f values into the equation equation

Nnesha (nnesha):